# The influence of spin in black hole triplets

### v3

@@ Abstract
Spin can influence the dynamics of the already chaotic black hole triplet system. We follow this problem in two sets of simulations: first, the Agekian-Anosova region (or region D), and second, using Pythagorean triangles. We use ARCcode, an N-body code that performs numerical integration of orbits. This code includes post-Newtonian corrections, which we include up to the 2.5th order. In set one of our simulations, we fix the masses of the black holes at 106 M. Then we run the simulations first without any spin added and after by initialising spin on one of the black holes. We find that after including spin into the system, 22% of the simulations changed outcomes. Either the systems went from having all black holes merging to having a black hole escaping the system, or vice versa. In the second set of simulations, we expanded into Pythagorean triangles as initial positions of black holes, stemming from Burrau’s three-body problem. We varied the masses of the black holes from 100 M - 1012 M. Black holes in these systems were given spin in normalised units ranging from 0 to $\sim$ 0.95. We find that black holes in the intermediary mass range of our simulations (104 M-107 M), were influenced the most by spin, particularly in their lifetimes. We also find that simulations, initialised as 2D, become 3D.
@@ 2.2 Initial Conditions
We focus first on the Agekian-Anosova homology region, also referred to as region D. This forms Set One of our simulations. This region describes the phase space region for zero initial velocity, equal mass three body systems [73, 72, 71]. In this space, two bodies are fixed at co-ordinates (0.5, 0) and (-0.5, 0) while the third body is placed in the region within the unit circle centered at (-0.5, 0) and bounded by the x and y axes. This is shown in Figure 1. By varying the position of M3 within this bounded region, a complete run of simulations is made.
@@
Table 1: A summary of Set One simulations. Black hole mass is fixed at 106 M. Placements of black holes are with respect to Figure 1. Each row represents a complete run of simulations.
Black Hole Position Spin Magnitude
M3 0
M3 0.95
-0.5 0.95
0.5 0.95
@@
We considered all of the Pythagorean triangles as an ensemble. We then focused more closely on the mass range of 104 M to 107 M with the Pythagorean triangles. This is summarised in Table 2.
Table 2: A summary of Set Two simulations. This is from the Pythagorean triangles. Each row represents a complete run of simulations.
Mass (M) Spin Vector
00, 101,…, 1011, 1012 [0 0 0], [0.05 0.05 0.05],…, [0.55 0.55 0.55]
104 , 104.2, …, 106.8, 107 As above
105.5 As above
2.3 Parameters
The main parameters we consider are: lifetimes of the systems, three-dimensional motion, evolution of some of the orbital elements and the fraction of mergers. Lifetime defines how long a system takes to evolve from three black holes into two. This can happen if two of the black holes merge (merger) or one of the black holes escapes the center of mass of the triplet (escape). The systems are always initialised in the 2D plane, but as we include spin this takes the system to 3D. We then follow the evolution of the system along the z-axis. For the orbital elements, we study this in Set Two. These are the orbital elements of one of the black holes with respect to the most massive, rotating black hole.
3 Results and Discussion
3.1 The Effect of Spin in Set One
The initial conditions were set so that two black holes (M1 and M2) of 1$\times$106 M each were fixed on the x-axis while the position of the third black hole (M3) of 10$\times$106 M was varied within the bounded region.This is described in Table 1 where 8094 individual simulations were conducted for each scenario: with spin and without spin.
Figure 3: The difference in log(lifetime/crossing time) between simulations in the Agekian Anosova region with and without spin on M3.
Figure 3 shows the difference in lifetimes between the cases with the inclusion of spin and without it. For this figure, the x and y axis show the initialised position of the black hole with spin, while the colour indicates the length of the lifetime for that particular setup. Here, a clearly defined region (the green portion) is seen. The addition of spin in the simulations can result in lifetimes of triplet systems increasing in some cases 30 magnitudes of order. This region of high variability corresponds to more non-hierarchical triplets. The effect of the additional element of rotation on M3 is likely due to the high sensitivity of these non-hierarchical triplets. Similar results are found when we analyse number of mergers and number of two-body encounters.
The outcomes of the black hole triplet system is also observed to change with the addition of spin in 22% of Set One simulations. The fraction of simulations that end via, merger, via escape, or do no change with the addition of spin, is shown in Figure 4.
3.2 The Effect of Increasing Spin
Here we focus on the ensemble of Pythagorean triangular configurations. In our previous study [3] we used an informal method of classifying the triangles into either more hierarchical or less hierarchical depending on the angle between the hypotenuse and the second longest side of the triangle. If there was an angle $>$ 30, the triangle was considered a non-hierarchical triangle and if the angle was $<$ 30, a hierarchical triangle. This is shown in Table 3.
@@ 3.2 The Effect of Increasing Spin
Triangle Classification Angle/
@@
When we consider the average lifetime of the triplet systems across all mass ranges ($10^{0}$ M - $10^{12}$ M) in Run Five, there is no discernible effect of spin above the mass 108 M. Mass becomes the dominant factor with large mass systems living for shorter times, which is expected. However, the intermediary mass range we observe variation in the lifetimes. In Figure 5, this mass range is looked at more closely from Run Six. It may be that spin has a small effect on the lifetimes around the 104 M - 105 M mass range. From this figure, a slight trend could be seen which we indicate with a black line. The equation of this line is, approximately, $y=2.25\times 10^{5}x-5.75\times 10^{4}$, where spin is on the x-axis and lg(Mass) is on the y-axis. It could also be that for these intermediate masses, the systems are just more sensitive and more chaotic, so we use this with caution. It is also obvious that there are other nearby regions that show high lifetimes but do not fall within the trend.
@@
Figure 5: The average lifetime for all triangular configurations for the mass range of 104 M - 107 M with increasing spin magnitude of the most massive black hole in each system. The black line indicates a potential trend, influenced by the increasing spin.
@@
One influence of spin is that it naturally transforms the 2D motion of the system into 3D. Initially, all bodies begin with co-ordinates limited to the xy-plane, however the spin vector of M1 is in the form [x y z]. We first consider the average among all triangles. There is an observable shift from the xy-plane to the xyz-plane upon the addition of spin to M1 illustrated in Figure 7, which shows the average maximum z co-ordinate achieved by the rotating black hole, M1, while the triplet system is intact (i.e. before any merger or escape). The added spin has the greatest effect on the motion in the z-axis at the large masses, while intermediary masses see some effect as well. This is interesting as the previously predictable large mass cases (such as 1010 M, 1011 M), obtain an added layer of complexity. However, the final outcomes such as lifetimes, mergers and the number of two-body encounters are not really affected by this as previously discussed.
@@
Figure 7: The lg(average ||||maximum z|||| co-ordinate/AU) of M1 throughout the lifetime of the triplet system as both mass and spin are varied for all triangular configurations.
@@
At the smaller mass units (100 M-102 M), the influence of spin is not very apparent but as the systems increase in mass, there is an increase in the maximum z-coordinate achieved by M1. This falls within the mass range of   104 M - 1010 M. As the mass unit increases beyond this, there is still 3-dimensional motion but much less than the intermediary masses.
@@
For Burrau’s 3,4,5 triangle which is considered chaotic - there is an observable effect of spin at intermediary and large mass units as seen in Figure 8. Smaller masses are not as influenced to move in the z-axis and larger masses end quickly before this type of motion can present itself.
@@
Figure 8: The lg(||||maximum z|||| co-ordinate/AU) of M1 throughout the lifetime of the triplet system as both mass and spin are varied for the 3,4,5 triangular configuration.
@@
Such behaviour is observed in the non-hierarchical triangles but more minimally in the hierarchical triangles at the smaller mass units. Interestingly, at larger mass units, the hierarchical triangles exhibit some z-motion before quick mergers take place as can be seen in Figure 9, where the results of the 13,84,85 triangle are shown. It is also more noticeable here that as the spin magnitude increases, the maximum z-coordinate also increases. The major difference is evident in how much the body M1 moves along the z-axis. For non-hierarchical triangles , the z value can become much larger than for the hierarchical triangles .
Figure 9: The lg(||||maximum z|||| co-ordinate/AU) of M1 throughout the lifetime of the triplet system as both mass and spin are varied for the 13,84,85 triangular configuration.
@@
These differences are typical as non-hierarchical configurations are thought to be more chaotic and unpredictable than the hierarchical ones [33, 34]. In all triangles at the extreme mass unit of 1012 M, systems end almost immediately without enough time to develop motion in the z-axis. These are the bands that appear in Figures 7, 8 and 9.
3.3 Effect on Orbital Elements
@@
We focus here on the evolution of the orbital elements of Set Two simulations. The argument of periapsis, $\omega$, and the longitude of ascending node, $\Omega$ are also affected by the added spin of M1. These orbital elements are shown in Figure 10. Periapsis is the point of closest approach in the orbit. The angle between the line of periapsis and where the orbital plane intersects the reference plane (light green and dark green respectively in Figure 10), is the argument of periapsis. This is shown in yellow in Figure 10. The longitude of the ascending node is the angle between the origin of the longitude of the reference plane (black) and the intersection of the orbital plane with the reference plane as the orbiting body ascends. This is shown in blue in Figure 10. We have observed that in some cases of this study, these two parameters fluctuate.
@@
As with the motion in the z-axis, these variations are mainly seen with the non-hierarchical triangles . In Figure 11, the panels on the left show the variation in $\omega$ and $\Omega$ for the 20,21,29 configuration and those on the right for 9,40,41 configuration. These represent the orbital elements of body M3 with respect to the rotating body, M1, for the mass unit of 1010 M in the 20,21,29 configuration and 1011 M for the 9,40,41. At these mass units there was maximum motion in the z-axis for the respective triangles. In black are the orbital elements over time when there is no spin in the system. The blue represents the orbital elements when M1 has maximal rotation of [0.55 0.55 0.55]. Here, it can be seen that the orbital elements express much more variation over time for the non-hierarchical triangles than the hierarchical triangles shown. This motion may be more complex for the non-hierarchical triangles due to longer lifetimes and more chaotic motion.
Figure 11: The change in $\omega$ (top) and $\Omega$ (bottom) for the 20,21,29 (left) configuration vs. the 9,40,41 configuration (right).
@@
Figure 12: The variation in $\omega$ as both mass unit of the system and spin magnitude of M1 are varied for the classical 3,4,5 configuration. On the left, the spin vector of M1 is [0.1 0.1 0.1] and on the right, [0.55 0.55 0.55]. The top, middle and bottom rows show, respectively, 100 M, 103 M and 106 M mass units.
@@
4 Astrophysical Implications
The motion of the system becomes more complicated as it moves from 2D to 3D and the orbital elements ($\omega$ and $\Omega$), show oscillatory motion in certain configurations more than others. This is of course, constrained to the distance unit of 1 parsec (206265 AU), which in larger mass cases, becomes more theoretical. Karttunen [64] showed that when initial velocities, were kept to zero, the motion was restricted to the plane. They derived statistical distributions from numerical orbital calculations of the binary energy within a triplet system for both the three-dimensional case and the planar case. They found that for each case the distributions were different statistically which could then affect, close encounters and lifetimes. Here, from our results (Figures 7), this is corroborated as the inclusion of spin of one of the black holes results in three dimensional motion. This of course, makes the systems more complex and their behaviour even more chaotic and unpredictable.
The lifetimes of the systems show some effect of added spin on the systems. At the intermediary mass range of $\approx$104 M - 107 M, there is a potential increase in lifetimes due to the rotation of M1.

### v4

@@ Abstract
Spin can influence the dynamics of the already chaotic black hole triplet system. We follow this problem in two sets of simulations: first, the Agekian-Anosova region (or region D), and second, using Pythagorean triangles. We use ARCcode, an N-body code that performs numerical integration of orbits. This code includes post-Newtonian corrections, which we include up to the 2.5th order. In set one of our simulations, we fix the masses of the black holes at 106 M. Then we run the simulations first without any spin added and after by initialising spin on one of the black holes. We find that after including spin into the system, 12.9% of the simulations changed outcomes. Either the systems went from having all black holes merging to having a black hole escaping the system, or vice versa. In the second set of simulations, we expanded into Pythagorean triangles as initial positions of black holes, stemming from Burrau’s three-body problem. We varied the masses of the black holes from 100 M - 1012 M. Black holes in these systems were given spin in normalised units ranging from 0 to $\sim$ 0.95. We find that intermediate mass black holes in the range of 104 M-105 M, were influenced the most by spin, particularly in their lifetimes. We also find that simulations, initialised as 2D, become 3D.
@@
[…]
[label7]organization=Visitor, Institute of Astronomy, University of Cambridge, Madingley Road,state=Cambridge, country=England
@@ 2.2 Initial Conditions
We focus first on the Agekian-Anosova homology region, also referred to as region D. This forms Set One of our simulations. This region describes the phase space region for zero initial velocity, equal mass three body systems [73, 72, 71]. In this space, two bodies are fixed at co-ordinates (0.5, 0) and (-0.5, 0) while the third body is placed in the region within the unit circle centered at (-0.5, 0) and bounded by the x and y axes. This is shown in Figure 1. By varying the position of M3 within this bounded region, a complete run of simulations is made. The bounded region was sampled by constructing 1000 circles centered at (-0.5,0) with decreasing radii. The upper limit for the radii was that of the initial unit circle and the lower limit was 0. Only quadrant one (x $>$ 0, y $>$ 0) was considered. Each of these circles were constructed of individual points with a stepsize of 0.01. These points for all of the circles created the set of positions in which the M3 body was placed during simulations. For each run, there were 8094 different positions of M3, making 8094 individual simulations per run.
@@
Table 1: A summary of Set One simulations. Following the setup shown in Figure 1. Each Run consists of 8094 individual simulations where M1 and M2 stay fixed on the negative and positive x-axis, respectively, but M3 moves each time within the bounded region. An initial run was conducted where none of the black holes had spin, as a control. To cover the Agekian-Anosova region fully, we also conduct runs three and four where the spinning black hole is fixed at M1 and M2, respectively. Black hole mass is fixed at 106 M, with the spinning black hole being more massive at 107 M for each case.
Run Black Hole Position Spin Magnitude Mass (106 M) (x,y) Position
One M1 0 1 (-0.5,0)
M2 0 1 (0.5,0)
M3 0 10 varying within region D
Two M1 0 1 (-0.5,0)
M2 0 1 (0.5,0)
M3 0.95 10 varying within region D
Three M1 0.95 10 (-0.5,0)
M2 0 1 (0.5,0)
M3 0 1 varying within region D
Four M1 0 1 (-0.5,0)
M2 0.95 10 (0.5,0)
M3 0 1 varying within region D
@@
In our previous study [3] we used an informal method of classifying the triangles into either more hierarchical or less hierarchical depending on the angle between the hypotenuse and the second longest side of the triangle. If there was an angle $>$ 30, the triangle was considered a non-hierarchical triangle and if the angle was $<$ 30, a hierarchical triangle. This is shown in Table 2.
@@
Triangle Classification Angle ()
@@
The effect of increasing spin was the focus for Set Two. We were interested in seeing how the effect of spin varied across the different masses of the black holes. To do this we constructed three different runs. In the first of run of set two simulations, we increased the mass of the black holes from 100 M to 1012 M in increasing factors of 10. For the second run we zoomed in more closely to the mass range of 104 M to 107 M in increasing factors of 100.2 M. For the third run we focused only on the 105.5 M mass since this mass was found to be the mass of our black hole triplets at which triplets preferred a merger outcome to an escape outcome [3]. For all of these runs, the most massive of the black holes, based on the Pythagorean triangles they were set up with. This spin varied from [0,0,0] to [0.55,0.55,0.55] increasing with [0.05,0.05,0.05]. These were kept along the diagonal for simplicity and not to introduce extra variability in an already complex system. The summary of the Set Two simulations is given in Table 3.
Table 3: A summary of Set Two simulations. This is from the Pythagorean triangles. For each of the Runs listed, all Pythagorean triangles with a hypotenuse $<$ 100 pc are used while the masses of the black holes are increased as listed.
Run log10(Mass (M)) Spin Vector
One 0, 1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11 [0 0 0], [0.05 0.05 0.05], [0.1, 0.1, 0.1], [0.15, 0.15, 0.15], [0.2, 0.2, 0.2], [0.25, 0.25, 0.25], [0.3, 0.3, 0.3], [0.35, 0.35, 0.35], [0.4, 0.4, 0.4], [0.45, 0.45, 0.45] [0.5, 0.5, 0.5], [0.55 0.55 0.55]
Two 4, 4.2, 4.4, 4.6, 4.8, 5, 5.2, 5.4, 5.6, 5.8, 6, 6.2, 6.4, 6.6, 6.8, 7 As above
Three 5.5 As above
2.3 Parameters
The main parameters we consider are: lifetimes of the systems, fraction of mergers, evolution of some of the orbital elements and how the motion is transformed from two-dimensional to three-dimensional. Lifetime defines how long a system takes to evolve from three black holes into two. This can happen if two of the black holes merge (merger) or one of the black holes escapes the center of mass of the triplet (escape). The systems are always initialised in the 2D plane, but as we include spin this takes the system to 3D. We then follow the evolution of the system along the z-axis. For the orbital elements, we study this in Set Two. These are the orbital elements of one of the black holes with respect to the most massive, rotating black hole.
3 Results and Discussion
3.1 The Effect of Spin in Set One
The initial conditions were set so that two black holes of 1$\times$106 M each were fixed on the x-axis while the position of the third black hole of 10$\times$106 M was varied within the bounded region.This is described in Table 1 where 8094 individual simulations were conducted for each scenario: with spin and without spin.
Figure 3: The difference in log10(lifetime/crossing time) between simulations in the Agekian Anosova region. This is the difference between Run One and Run Two from Table 1. Here the masses of the fixed black holes are 106 M. The spinning black hole varies within the region D in every simulation and has a mass of 10$\times$106 M.
Figure 3 shows the difference in lifetimes between the cases with the inclusion of spin and without it. For this figure, the x and y axis show the initialised position of the black hole with spin, while the colour indicates the length of the lifetime for that particular setup. Here, a clearly defined region (the green portion) is seen. The addition of spin in the simulations can result in lifetimes of triplet systems increasing in some cases 30 magnitudes of order. This region of high variability corresponds to more non-hierarchical triplets. In these non-hierarchical triplets, there are initial configurations that are very similar but live magnitudes longer. This corresponds to the red dots within the green region. The chaoticity of the three-body problem is most likely the cause of this. Very small changes in the initial conditions result in very large difference in outcomes. In some cases here, almost identical setups have a lifetime difference of 1020 years. The effect of the additional element of rotation on M3 is likely due to the high sensitivity of these non-hierarchical triplets. Similar results are found when we analyse number of mergers and number of two-body encounters.
The outcomes of the black hole triplet system is also observed to change with the addition of spin in 12.9% of Set One simulations. The fraction of simulations that end via, merger, via escape, or do no change with the addition of spin, is shown in Figure 4.
3.2 The Effect of Increasing Spin in Set Two
@@
Here we focus on the ensemble of Pythagorean triangular configurations (Set Two). When we consider the average lifetime of the triplet systems across all mass ranges ($10^{0}$ M - $10^{12}$ M) in Set Two Run One (Refer to Table 3), there is no discernible effect of spin above the mass 108 M. Mass becomes the dominant factor with large mass systems living for shorter times, which is expected. However, the intermediary mass range we observe variation in the lifetimes. In Figure 5, this mass range is looked at more closely from Set Two Run Two. It may be that spin has a small effect on the lifetimes around the 104 M - 105 M mass range.
@@
Figure 5: The average lifetime, in crossing time, for all Pythagorean triangular configurations within the mass range of 104 M - 107 M and with increasing spin magnitude.
@@
One influence of spin is that it naturally transforms the 2D motion of the system into 3D. Initially, all bodies begin with co-ordinates limited to the xy-plane, however the spin vector of the rotating black hole is in the form [x y z]. We first consider the average among all triangles. There is an observable shift from the xy-plane to the xyz-plane upon the addition of spin illustrated in Figure 7, which shows the average maximum z co-ordinate achieved by the rotating black hole, while the triplet system is intact (i.e. before any merger or escape). This z-motion is normalised to the hypotenuse for each triangle. The added spin has the greatest effect on the motion in the z-axis at the large masses, while intermediary masses see some effect as well. This is interesting as the previously predictable large mass cases (such as 1010 M, 1011 M), obtain an added layer of complexity. However, the final outcomes such as lifetimes, mergers and the number of two-body encounters are not really affected by this as previously discussed.
@@
Figure 7: The log10(average ||||maximum z|||| co-ordinate/AU) of the rotating black hole throughout the lifetime of the triplet system as both mass and spin are varied for all triangular configurations.
@@
At the smaller mass units (100 M-102 M), the influence of spin is not very apparent but as the systems increase in mass, there is an increase in the maximum z-coordinate achieved by the rotating black hole. This falls within the mass range of   104 M - 1010 M. As the mass unit increases beyond this, there is still 3-dimensional motion but much less than the intermediary masses.
@@
For Burrau’s 3,4,5 triangle which is considered chaotic - there is an observable effect of spin at intermediary and large mass units as seen in Figure 8 (a). Smaller masses are not as influenced to move in the z-axis and larger masses end quickly before this type of motion can present itself.
(a) 3,4,5 triangle
(b) 13,84,85 triangle
@@
Figure 8: The lg(||||maximum z|||| co-ordinate/AU) of the spinning black hole throughout the lifetime of the triplet system as both mass and spin are varied for the 3,4,5 triangular configuration (a) and the 13,84,85 triangular configuration (b).
@@
Such behaviour is observed in the non-hierarchical triangles but more minimally in the hierarchical triangles at the smaller mass units. Interestingly, at larger mass units, the hierarchical triangles exhibit some z-motion before quick mergers take place as can be seen in Figure 8 (b), where the results of the 13,84,85 triangle are shown. It is also more noticeable here that as the spin magnitude increases, the maximum z-coordinate also increases. The major difference is evident in how much the body the rotating black hole moves along the z-axis. For non-hierarchical triangles , the z value can become much larger than for the hierarchical triangles .
@@
These differences are typical as non-hierarchical configurations are thought to be more chaotic and unpredictable than the hierarchical ones [33, 34]. In all triangles at the extreme mass unit of 1012 M, systems end almost immediately without enough time to develop motion in the z-axis. These are the bands that appear in Figures 7 and 8.
@@
From the Set One simulations, we also observe an effect on the orbital elements of our systems. The argument of periapsis, $\omega$, and the longitude of ascending node, $\Omega$ are affected by the added spin in the system. These orbital elements are shown in Figure 9. Periapsis is the point of closest approach in the orbit. The angle between the line of periapsis and where the orbital plane intersects the reference plane (light green and dark green respectively in Figure 9), is the argument of periapsis. This is shown in yellow in Figure 9. The longitude of the ascending node is the angle between the origin of the longitude of the reference plane (black) and the intersection of the orbital plane with the reference plane as the orbiting body ascends. This is shown in blue in Figure 9. We have observed that in some cases of this study, these two parameters fluctuate.
@@
As with the motion in the z-axis, these variations are mainly seen with the non-hierarchical triangles . In Figure 10, the panels on the left show the variation in $\omega$ and $\Omega$ for the 20,21,29 configuration and those on the right for 9,40,41 configuration. These represent the orbital elements of black hole M3 (non-rotating) with respect to M1 (rotating) for intermediary the mass unit of 104.5 M, the mass at which we see the most influence of spin in the lifetime of our systems. In blue are the orbital elements over time when there is no spin in the system. The orange represents the orbital elements when M1 has maximal rotation of [0.55 0.55 0.55]. Here, it can be seen that the orbital elements express much more variation over time for the non-hierarchical triangles than the hierarchical triangles shown. This motion may be more complex for the non-hierarchical triangles due to longer lifetimes and more chaotic motion. For the simulations without spin, there is still slight variation in periapsis which we expect from relativistic precession of the periapsis.
(a) 20,21,29 triangle
(b) 9,40,41 triangle
(c) 20,21,29 triangle
(d) 9,40,41 triangle
Figure 10: The change in $\omega$, the argument of periapsis, (top) and $\Omega$, the longitude of ascending node, (bottom) for the 20,21,29 (left) configuration vs. the 9,40,41 configuration (right). In all cases the masses of the black holes are maintained at 104.5 M within each triangular configuration.
@@
Figure 11: The variation in $\omega$ as both mass unit of the system and spin magnitude of M1 (rotating black hole) are varied for the classical 3,4,5 configuration. On the left, the spin vector of M1 is [0.1 0.1 0.1] and on the right, [0.55 0.55 0.55]. The top, middle and bottom rows show, respectively, 100 M, 103 M and 106 M mass units.
@@
4 Discussion and Conclusion
[…]
In Set Two, we have found that while there is not a significant effect of spin on some parameters like the fraction of mergers, the lifetimes of these systems may have been affected, particularly for intermediate mass black holes in the range 104 M-105 M.
We have also found that certain effects are more visible in the non-hierarchical triangles than in the hierarchical triangles like the motion of the orbital elements $\omega$, the argument of periapsis, and $\Omega$, the longitude of ascending node, owing to the chaotic tendencies of the non-hierarchical configurations. The inclusion of spin which is not perpendicular to the orbital plane leads to 3D motion. This is of course, constrained to the distance unit of 1 parsec (206265 AU), which in larger mass cases, becomes more theoretical.
Karttunen [64] showed that when initial velocities were kept to zero, the motion of triplets was restricted to the plane. They derived statistical distributions from numerical orbital calculations of the binary energy within a triplet system for both the three-dimensional case and the planar case. They found that for each case the distributions were different statistically which could then affect, close encounters and lifetimes. Here, from our results (Figures 7), this is corroborated as the inclusion of spin of one of the black holes results in three dimensional motion. This of course, makes the systems more complex and their behaviour even more chaotic and unpredictable.